What is the Graph of y = |x|.(x-2) ? Plz answer it someone

Grade:12

2 Answers

Vikas TU
5 years ago
Dear Student,
for x less than 0 , y= x(x-2) = x^2 – 2x
        For x greater than 0 , y= -x(x-2) = -x^2 + 2x 
These are the two functions for 2 above intervals.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Parag agrawal
5 years ago
For X > 0 x^2 - 2x we get an upward parabola with zeroes at X = 0,2But the graph will not extend towards negative values of x ( as we took x to be greater than zero)For X